OpenStudy (anonymous):

In 5 minuets, Peter completed 2 math problems. Yvonne says he did 3/5 of a problem each minuets. Is she correct? Explain!

OpenStudy (anonymous):

This is a rate problem. A rate problem is a fraction of two variables. And rates are fixed so. \[\frac{problems}{minutes}=\frac{2 problems}{5 minutes}=\frac{3/5}{1}\] If true, then the rate is maintained and she's right. So it comes down to, are these equal: \[\frac{ 2}{5}=\frac{ 3/5}{1}\] or\[\frac{ 2}{5}=\frac{ 3}{5}\]

OpenStudy (anonymous):

so she is right?

OpenStudy (anonymous):

Well if \[\frac{2}{5}=\frac{3}{5}\] then yes. If they are not equal, then no.

OpenStudy (anonymous):

ok well thank you so much i didnt know what to tell my cousin or how to help her